### v1

# A Foundation Model for General Moving Object Segmentation in Medical Images

Abstract
Medical image segmentation aims to delineate the anatomical or pathological structures of interest, playing a crucial role in clinical diagnosis. A substantial amount of high-quality annotated data is crucial for constructing high-precision deep segmentation models. However, medical annotation is highly cumbersome and time-consuming, especially for medical videos or 3D volumes, due to the huge labeling space and poor inter-frame consistency. Recently, a fundamental task named Moving Object Segmentation (MOS) has made significant advancements in natural images. Its objective is to delineate moving objects from the background within image sequences, requiring only minimal annotations. Thus, it holds the potential to mitigate the annotation workload. In this paper, we propose the first foundation model, named iMOS, for MOS in medical images. Extensive experiments on a large multi-modal medical dataset validate the effectiveness of the proposed iMOS. Specifically, with the annotation of only a small number of images in the sequence, iMOS can achieve satisfactory tracking and segmentation performance of moving objects throughout the entire sequence in bi-directions. We hope that the proposed iMOS can help accelerate the annotation speed of experts, and boost the development of medical foundation models.
Index Terms—  Medical image, Moving object segmentation, Foundation model
1 Introduction
The accurate segmentation of specific regions in medical images is crucial for clinical diagnosis and treatment planning. Recently, deep learning-based segmentation algorithms have shown promising results in medical images [1, 2, 3]. Most of them focused on the segmentation of only one single modality or specific anatomical structure. Different from them, the Segment Anything Model (SAM) [4] was proposed to achieve general image segmentation, and lots of its extension works analyzed the medical SAM [5, 6, 7]. In their methods, using a few manual prompts (i.e., points or boxes) can guide the SAM to achieve impressive segmentation accuracy on specific medical images. However, directly applying these 2D models to videos or 3D volumes often presents additional challenges [8]. Moreover, they required substantial manual prompts for segmenting multiple frames/slices or 3D volumes. This limits their flexibility and convenience in real-life usage scenarios.
Most recently, several studies have been proposed to explore the novel task, named Moving Object Segmentation (MOS). It aims to automatically track and segment moving objects in consecutive frames with minimal or no annotations (i.e., semi-supervised and unsupervised MOS). In this study, we focus on exploring the semi-supervised MOS (see Figure 1), due to the structure complexity and scene variability of medical data. Compared to the unsupervised MOS, the semi-supervised one can ensure the accuracy of segmentation while maintaining flexibility. Semi-supervised MOS has already demonstrated impressive performance in natural images [9, 10, 11]. However, there is still a lack of research that investigates MOS in medical images currently.
Fig. 1: Pipeline of semi-supervised MOS. The leftmost image represents the first frame with manual annotation, while the other images are the segmentation results generated by semi-supervised MOS method.
In this work, we proposed a foundation model for general MOS in medical images named iMOS. To the best of our knowledge, we are the first to explore MOS in medical images. Additionally, we validated the proposed method on a multi-modal and large-scale medical dataset. The experimental results indicate that with minimal annotations, iMOS can effectively track and segment anatomical structures in medical videos and volumes. By labeling one single frame, iMOS can achieve forward and backward segmentation in the sequence, even for those objects unseen during training. This proves the strong generalization of our proposed foundation MOS model.
2 Related Work
Current MOS can primarily be categorized into unsupervised and semi-supervised approaches. Different from general deep learning, the supervision here only refers to manual interaction during testing rather than the training [12].
Unsupervised MOS. It aims to segment salient objects sequences without any human intervention. Traditional methods used similarity calculation [13], motion boundaries [14], and thresholds [15], etc., to segment objects. These methods are easily affected by occlusion, contrast, and complex structures. Deep-learning-based methods can better model object motion information using techniques, including optical flow [16], temporal coherence [17], and motion information enhancement [18]. However, due to the lack of supervision, the above methods are poor in robustness and effectiveness.
Semi-supervised MOS. In this kind of method, the masks of the first frame or the first few frames are first given for supervision. Then, the model segments subsequent frames based on the cues of these frames. This method performs well in various computer vision tasks. Representative approaches rely on data augmentation [19], yet their effectiveness is limited, and they often suffer from slow processing speeds. Tracking-based methods [20] are proficient at monitoring moving objects. However, they tend to accumulate errors over time, resulting in diminished robustness. At present, memory-based methods [9, 10, 11] have emerged as the primary approach for MOS, primarily owing to their outstanding performance and straightforward architecture. However, the above methods are all constructed for natural image segmentation. Hence, it is highly desired for the establishment of a foundational medical MOS model.
3 Methodology
3.1 Semi-supervised MOS
We consider the segmentation of anatomical structures in medical images, whether in videos or 3D volumes, as a semi-supervised MOS task. Specifically, we consider frames in video or slices in 3D image which contain the anatomical structures of interest as a image sequence. Additionally, manual annotations for the first frame are provided, and the segmentation of subsequent images is completed by MOS method. Figure 2 illustrates the architecture of our method which is based on XMem [11] model, it only requires manual annotation of a single frame to complete the segmentation for all frames.
Fig. 2: Architecture of iMOS. Lines with different colors are pathways of different memory systems.
3.1.1 Memory Reading
iMOS consists of three trainable modules: query encoder, decoder, value encoder, and three memory systems: sensory memory, working memory, long-term memory. For a given sequence of images and the mask of the target objects in the first frame, sensory memory and working memory are initialized with the first frame and its corresponding mask. For the subsequent images, the readout features $F$ are extracted from the memory value $v$ based on the affinity $W$ between query $q$ and memory key $k$. Simultaneously, a mask is generated by the decoder with $F$ and the hidden representation, $h_{t-1}$, from the previous time step in the sensory memory. Additionally, the value encoder encodes the image and mask every $r$-th frames to generate new memory value. The aforementioned memory reading process can be expressed as: $F=vW(k,q)$.
3.1.2 Memory Updating
The memory systems in iMOS update at various frequencies to store features at different time scales. Sensory memory updates at the fastest frequency to retain short-term information, utilizing multi-scale features from the decoder at every frame and updating through gated recurrent units (GRU). Furthermore, at every $r$-th frame, the memory value generated by the value encoder are used to update sensory memory via another GRU, enabling sensory memory to discard redundant information. Working memory is employed for precise matching over a few seconds, where memory value and memory key copied from query are added to the working memory every $r$-th frames. To prevent working memory from occupying excessive memory space, the number of frames in working memory is restricted between $T_{min}$ and $T_{max}$. Additionally, long-term memory is introduced to store compact and representative features. When working memory reaches the upper limit of its storage capacity, $T_{max}$, all frames except the first frame and the most recent $T_{min}-1$ frames are added to long-term memory through a prototype selection and memory potentiation algorithm.
Prototype Selection. Prototype selection aims to choose a representative subset, denoted as prototype keys $k^{p}$, from the candidate keys $k^{c}$. These subsets have the highest usage frequency within working memory, which is defined by the cumulative total affinity in the affinity matrix $W$.
Memory Potentiation. The prototype selection results in sparse prototype keys. If the same method were applied to extract prototype values, it might not adequately represent other candidates. Memory potentiation is introduced to address this issue. Specifically, similar to memory reading, candidate value are extracted with affinity matrix $W(k^{c},k^{p})$ between candidate key $k^{c}$ and prototype key $k^{p}$: $v^{p}=v^{c}W(k^{c},k^{p})$
3.2 Parameter Efficient Tuning
The training process of iMOS is divided into training on static images, training on long sequences, and training on short sequences. Training the model from scratch or fine-tuning the entire model is time-consuming. To facilitate rapid and efficient fine-tuning of the model on large-scale dataset, we introduced the adapter module based on the Conv. Parallel structure mentioned in [21], which achieves the best trade-off between parameter efficiency and performance. We froze all parameters of the query encoder and value encoder, then we applied the adapter on them, enabling them to learn residual representations between medical images and natural images. The architecture of adapter and our implementation scheme are illustrated in Figure 3 and Figure 4, respectively.
Fig. 3: Architecture of adapter module. It comprises a convolutional layer, a transposed convolutional layer, batch normalization, ReLU activation, and a learnable parameter, $\alpha$.
Fig. 4: Implementation scheme of adapter. Left: bottleneck with adapter module in key encoder. Right: basicblock with adapter module in value encoder.
4 Experimental Results
4.1 Dataset and Training Details
To objectively evaluate the performance of the proposed method on the medical images, we collected data from multiple public datasets [22, 23, 24], including four modalities: CT, MRI, endoscopy, and electron microscopy (see Figure 5). The collected dataset encompassed various categories, including but not limited to the liver, kidneys, gallbladder, stomach, adipose tissue, surgical instruments (e.g., L-shaped hook electrocautery), and cellular components (e.g., mitochondria).
Fig. 5: Examples of the four modalities in the dataset.
For each category of video and 3D volume, we construct a dataset in which objects are tracked from the first frame to the last frame. For image sequences generated from CT and MRI, we excluded categories with short length. The basic information of the dataset is presented in Table 1.
Table 1: Dataset description.
  Modality Source Category Video/3D volume Generated sequence
CT AMOS22 [22] 7 300 2091
MRI 7 60 419
Endoscopy CholecSeg8k [23] 12 101 597
Electron Microscopy EPFL [24] 1 2 109
We split the video or 3D volume into a 3:1:1 ratio for training/validation/testing. We selected only a subset of categories for training to evaluate generalization across different categories. The optimal model was chosen based on the results from the validation set. All images were padded to have equal width and height before resizing to a resolution of $480\times 480$. Additionally, given the wide CT value range in CT data, we adjust appropriate window widths and window levels according to different anatomical regions to exclude irrelevant information and reduce interference. All other experimental settings remained consistent with those in XMem.
4.2 Evaluation
Based on our proposed framework and the problem it solves, we design two experiments. The first is the validity experiment. We compared the performance of the model before and after fine-tuning on medical image segmentation. The second is a stability experiment. We use the fine-tuned model to segment multiple organs in multiple modalities to prove that iMOS has good performance in multi-organ segmentation.
Table 2: Validity experiment
[…]
MRI 0.626(0.292) 0.542(0.311) 0.710(0.291) 0.806(0.225) 0.759(0.256) 0.854(0.208)
Endoscope 0.850(0.162) 0.853(0.142) 0.852(0.140) 0.856(0.138) 0.854(0.160) 0.858(0.140)
[…]
Table 3: Stability experiment. Categories marked with an asterisk ($*{}^{*}$) indicate that the class did not appear in the training set.
[…]
CT Spleen 0.893(0.168) 0.855(0.196) 0.932(0.160) Endoscope Abdominal Wall 0.886(0.057) 0.928(0.085) 0.844(0.110)
Right kidney 0.909(0.133) 0.871(0.165) 0.948(0.126) Liver 0.895(0.054) 0.910(0.087) 0.880(0.103)
Left kidney 0.904(0.158) 0.862(0.191) 0.945(0.134) Gastrointestinal Tract 0.816(0.126) 0.811(0.170) 0.822(0.153)
$*{}^{*}$ 0.810(0.200) 0.683(0.214) 0.937(0.143) Fat 0.872(0.062) 0.906(0.095) 0.837(0.097)
$*{}^{*}$ 0.782(0.270) 0.760(0.395) 0.804(0.258) Grasper 0.818(0.162) 0.750(0.198) 0.887(0.149)
$*{}^{*}$ 0.780(0.208) 0.752(0.243) 0.809(0.231) Gallbladder 0.876(0.161) 0.875(0.142) 0.861(0.171)
$*{}^{*}$ 0.863(0.169) 0.800(0.227) 0.945(0.140) Connective $*{}^{*}$ 0.837(0.065) 0.858(0.095) 0.816(0.146)
MRI Spleen 0.921(0.091) 0.885(0.126) 0.956(0.091) $*{}^{*}$ 0.636(0.205) 0.583(0.220) 0.689(0.182)
Right kidney 0.917(0.078) 0.893(0.104) 0.940(0.083) Cystic $*{}^{*}$ 0.870(0.040) 0.806(0.065) 0.935(0.068)
Left kidney 0.907(0.100) 0.877(0.131) 0.937(0.106) L-hook $*{}^{*}$ 0.912(0.080) 0.906(0.099) 0.918(0.106)
$*{}^{*}$ 0.686(0.245) 0.513(0.249) 0.859(0.209) Hepatic $*{}^{*}$ 0.729(0.186) 0.556(0.207) 0.901(0.098)
$*{}^{*}$ 0.775(0.215) 0.756(0.263) 0.795(0.219) Liver $*{}^{*}$ 0.945(0.004) 0.973(0.007) 0.918(0.043)
$*{}^{*}$ 0.727(0.240) 0.662(0.263) 0.793(0.216) Electron Microscopy Mitochondria 0.879(0.144) 0.805(0.149) 0.954(0.143)
$*{}^{*}$ 0.725(0.289) 0.649(0.311) 0.802(0.281)
We use quantitative evaluation indicators commonly used in general image segmentation tasks, namely regional similarity $\mathcal{J}$ [25]and contour similarity $\mathcal{F}$ [26]. Among them, $\mathcal{J}$ is a measure of regional accuracy by calculating IoU. $\mathcal{F}$ is still based on IoU, but it only focuses on the coincidence of edge areas. In addition, we use the average value of $\mathcal{J}$ and $\mathcal{F}$ as an indicator named $\mathcal{J}$$&\mathcal{F}$, which is also widely used to measure the performance of MOS models.
Table 2 summarizes the segmentation results before and after iMOS fine-tuning in 4 modalities (CT, MRI, endoscopy and electron microscopy). It can be found that under the three quantitative evaluation indicators, the fine-tuned model has an absolute leading advantage in medical image segmentation.
Table 3 shows the segmentation capabilities of iMOS when facing organs of different modalities. Among them, the CT and MRI data sets each contain 7 types of organs, the endoscope contains 12 types of segmentation objects, and the electron microscope contains one type. The organs marked (*) represent that the data of this organ did not appear during the training of iMOS. iMOS performs well in segmentation on these medical images.
Fig. 6: Examples of inference results of multiple modalities. In each row, the leftmost frame is the image with manual annotations, and the segmentation results for the remaining frames are obtained through inference from this frame.
Fig. 7: Qualitative results of 3D CT volume segmentation of lung organs. The first row of XY planes is the transverse plane segmentation map, the second row of XZ planes is the coronal plane segmentation map, and the third row of YZ planes is the sagittal plane segmentation map.
Figure 6 comprehensively shows the qualitative results of iMOS under CT, MRI, endoscopy and electron microscopy. With the help of the mask prompted by the first frame, iMOS can track and segment targets from multiple organs in multiple modalities with good performance.
Figure 7 visualizes the experimental results of 3D CT lung organ segmentation. We provided three cross-sectional views: neutral plane, transverse plane, and sagittal plane. The obtained 3D segmentation results are obtained based on the first frame of transverse plane.
Fig. 8: Bi-directional segmentation results. the segmentation organ in the CT image is kidney. iMOS can not only segment backward according to the prompts of the first frame, but also segment forward according to the prompts of the last frame.
Figure 8 illustrates the experimental results of bi-directional image segmentation. The segmentation conditions of iMOS are not limited to the first frame prompt. It can also perform reverse inference based on the segmentation hints of the last frame or intermediate frames.Forward and backward predictions have similar effects
5 Discussion and Conclusion
5.1 Discussion
The proposed foundation model iMOS provides technical support for video segmentation related tasks. Here are some examples of practical applications.
Medical Image Annotation. iMOS can automatically segment the targets of other frames of the video based on the mask prompt of the first frame of the video. iMOS can also regard the 3D volume as a multi-frame video, so medical experts can adjust the segmentation results given by iMOS from videos or 3D volumes based on the first frame prompt. In addition, the segmentation results generated by iMOS have front-to-back consistency, which can avoid inconsistent mask positions in the front and back frames of the video due to human judgment bias. Based on the above points, iMOS can improve the efficiency and accuracy of medical image annotation.
Surgical Navigation. iMOS can process multiframe videos and continuously track the expected target in long videos based on the mask of the first frame. Therefore, it has advantages in surgical navigation.
iMOS is able to achieve better tracking and segmentation effects in most scenes. However, some organs or structures are not suitable for segmentation using iMOS. For example, if the structural gap between different organs is small, iMOS may track other organs during segmentation. For organs with separated structures, such as blood vessels, iMOS may ignore the structure of the separated parts.
5.2 Conclusion
In this work, we propose iMOS, a foundation model for general moving object segmentation in medical images. iMOS can automatically segment the targets of the remaining frames of the video based on the segmentation mask of the first frame provided by the video. At the same time, we regard each slice of medical 3D volume as each frame of the video, so iMOS can also perform segmentation on medical 3D volume. We have demonstrated its validity and stability on medical images through a series of experiments. iMOS provides technical accumulation for auxiliary medical video or 3D volume annotation tasks and surgical navigation.